Unit 2: Something Is Going On – Visual Diagrams Of Daily Changes
Both NISA and the IAEA think there is enough of a chance that something is going on inside unit 2 that they have asked TEPCO to create alternative ways of measuring temperatures. IAEA also mentioned that the manner in which sensors fail does not match some of the behaviors seen on sensors in unit 2, hinting that sensors labeled failed may not be.
In order to help understand what is going on inside unit 2 we have developed some visual diagrams of daily readings that bring together data from various documents. The position along the horizontal plane for each sensor is labeled with a degree if it is known. Others are marked as close as possible to the structure or known landmark for location of that sensor. Water rates, pressure, and hydrogen concentrations are also included.
Click on each image for a larger view. If these prove useful we will continue them for a while. There is also much valuable information in our reactor graphs page that show the entire historical trends of each reactor’s sensors. There are also some notes on trends at unit 2 at the bottom of this article.
Below are some notes on changes noticed recently. A number of sensors have changed since February 19th. What factors lead to this still needs to be clearly determined.
Unit 2 continues to show unusual readings. TEPCO insists it is all faulty gauges and media hype.
The CRD Housing is reading -5.8c (spiked very high an dropped very low quickly)
RPV Bellows 137.8c (down but still high)
Drywell HVH 54.3c (up steady)
The other two RPV bottom head sensors have been heating up since the 19th of February. The sensors at 135° and 270° were both at 29c and have now risen to 45.8c and 39c The support skirt junctions at 135° and 270° have also been on the rise since around February 19th.
The Safety Relief Valve has also climbed upward and is currently at 51.7c. It has also been on the increase since February 19th. This valve is high up and away from the RPV in the steam line.
Radiation in the Torus B sensor began jumping upward about January 22nd. It has then been bouncing between 5-20 Sv/h ever since. TEPCO labels it bad but the sensor was on a steady downward path before January 22nd.
Pressure in unit 2 began to increase faster on February 19th. It peaked about 120 kPa and dropped to 118 kPa today.
Core Spray System: Approx.6.0 m3/h
Feed Water System: Approx. 2.9 m3/h
This article would not be possible without the extensive efforts of the SimplyInfo research team
Join the conversation at chat.simplyinfo.org
© 2011-2023 SimplyInfo.org, Fukuleaks.org All Rights Reserved Content cited, quoted etc. from other sources is under the respective rights of that content owner. If you are viewing this page on any website other than http://www.simplyinfo.org (or http://www.fukuleaks.org) it may be plagiarized, please let us know. If you wish to reproduce any of our content in full or in more than a phrase or quote, please contact us first to obtain permission.